how will you prepare 50mills of 50 millimoles solution of potassium hydroxide and calcium hydroxide given that their molecular weight are 56.1g/mol and 74.0g/mol respectively
1
Expert's answer
2021-08-31T01:31:58-0400

CM = n/V

n = m/M


1) 0.05 M = n/0.05 L

n (KOH) = 0.05/0.05 = 1 mol

m (KOH) = n x M = 1 x 56.1 = 56.1 g

To prepare 50mills of 50 millimoles solution of potassium hydroxide, 56.1 g of KOH is required.


2) 0.05 M = n/0.05 L

n (Ca(OH)2) = 0.05/0.05 = 1 mol

m (Ca(OH)2) = n x M = 1 x 74.0 = 74.0 g

To prepare 50mills of 50 millimoles solution of potassium hydroxide, 74.0 g of Ca(OH)2 is required.